OpenStudy (samirahdanyel):

The axis of symmetry for the graph of the function is f(x) =1/4 x2 + bx + 10 is x = 6. What is the value of b?

OpenStudy (rayep12):

is that what you mean?

OpenStudy (samirahdanyel):

yes

OpenStudy (rayep12):

\[\frac{ 1 }{ 4 }x ^{2}\] would become \[\frac{ 1 }{ 4 }*6^{2}\]

OpenStudy (rayep12):

so \[6^{2}\] is what?

OpenStudy (samirahdanyel):

36

OpenStudy (rayep12):

great

satellite73 (satellite73):

not to butt in, but that is not what the question is asking the axis of symmetry of \[y=ax^2+bc+c\] is \[x=-\frac{b}{2a}\]

satellite73 (satellite73):

in your case \(a=\frac{1}{4}\) `

OpenStudy (rayep12):

ohh okay u help him sate i got someone to help

satellite73 (satellite73):

set \[-\frac{b}{2\times \frac{1}{4}}=6\]and solve for \(b\)

OpenStudy (samirahdanyel):

okay gimmie a minute

OpenStudy (samirahdanyel):

i got -3

satellite73 (satellite73):

seems right to me

satellite73 (satellite73):

wanna check?

OpenStudy (samirahdanyel):

sure

satellite73 (satellite73):

http://www.wolframalpha.com/input/?i=vertex+of++x%5E2%2F4-3x%2B10 since the x coordinate of the vertex is 6, you are right

OpenStudy (samirahdanyel):

okay thank you

satellite73 (satellite73):

yw